# natural language processing; greenwashing detection; corporate climate-related greenwashing; climate pretrained models; climate topic detection; green claim detection; climate QA; climate-related deceptive techniques; enviromental performance prediction copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processing

### v1

@@
Abstract.
Greenwashing is an effort to mislead the public about the carbon transition of an entity, such as a state or company. We provide a comprehensive survey of the scientific literature addressing natural language processing methods to identify potentially misleading climate-related corporate communications, indicative of greenwashing. We break the detection of greenwashing into intermediate tasks, and review the state-of-the-art approaches for each of them. We discuss datasets, methods, and results, as well as limitations and open challenges. We also provide an overview of how far the field has come as a whole, and point out future research directions.
natural language processing; greenwashing detection; corporate climate-related greenwashing; climate pretrained models; climate topic detection; green claim detection; climate QA; climate-related deceptive techniques; enviromental performance prediction
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processing
1. Introduction
@@
Since the first report of the Intergovernmental Panel on Climate Change in 1990  (Change, 1990), world leaders and citizens have increasingly recognized the significant adverse impacts of human activities on the environment, particularly on climate change (Peter Jackson, 2007). This increased awareness has translated into guidelines, laws, and investments, such as the European Green Deal (Publications Office of the European Union, 2019) or the Inflation Reduction Act in the US (The White House, 2023). Many companies have used the financial incentives offered by states, and the guidelines and legislation to make significant steps towards sustainability (Time, 2024). At the same time, this growing attention also generated an advertising opportunity for companies that aim to promote themselves as environmentally aware and responsible. Indeed, some companies have been found to deliberately manipulate their data and statistics to appear more environment-friendly. The Diesel Scandal around the Volkswagen car company is a prominent example (US Environmental Protection Agency, 2024). However, such cases are not the norm. More commonly, companies avoid outright data manipulation but present themselves in a misleadingly positive light regarding their environmental impact – a practice called greenwashing.
In addition to their broader communication strategies, most companies disseminate information about their environmental initiatives through dedicated sustainability 10-K and annual reports. In addition, corporations also leverage various media channels and traditional communication platforms to convey their environmental credentials. As a result, the corpus of documents used in textual analysis is diverse and heterogeneous.
In such documents, greenwashing can appear in multiple forms. The Green Claims Directive of the European Union (US Environmental Protection Agency, 2023) identifies four types of greenwashing:
  1. (1)
    using unreliable sustainability labels (e.g., H&M and Decathlon have used labels such as eco design and conscious and were prompted by the Netherlands Authority for Consumers and Markets to stop the practice in 2022 as the retailers did not clearly describe why the products received those labels (Netherlands Authority for Consumers and Markets (2022), ACM))
  2. (2)
    presenting legal requirements for the product as its distinctive features (e.g., McDonald’s advertises its reduction of plastic waste by using reusable cutlery (McDonald’s France, 2024), while this is a legal obligation since January 2023 in France (LegiFrance, 2020))
  3. (3)
    making green claims about the entire product when the claim applies only to a part/aspect of the product (e.g., Kohl’s and Walmart were fined by the U.S. Federal Trade Commission for the misleading claim that their products were made of bamboo, a sustainable material, when they were made out of rayon, which is derived from bamboo via a highly toxic process that removes any environmental benefits of bamboo (Lesley Fair, Federal Trade Commission, 2022)), and finally
  4. (4)
    making environmental claims for which the company cannot provide evidence (e.g., Keurig was fined by Canada’s Competition Bureau for misleading claims about the recycling process of its single-use coffee pods (Competition Bureau Canada, 2022)).
@@
Greenwashing is a widely observed practice rather than a marginal occurrence: the Australian Competition & Consumer Commission performed a sweep analysis on 247 businesses in October 2022 and found that $57\%$ of them made misleading claims about their environmental performance (Australian Competition and Consumer Commission, 2023). The International Consumer Protection Enforcement Network, too, found in November 2020 that $40\%$ of around 500 websites of businesses covering different sectors contain misleading green claims (UK Competition and Markets Authority, 2021).
@@
Addressing greenwashing is not only in the broader public interest but also in the company’s best interest, as the practice is often subject to legal sanctions prohibiting misleading or deceptive conduct, with businesses facing fines or lawsuits as a consequence (Mary Foley, Forbes, 2024). In addition, political awareness and public awareness of the issue via traditional and social media (Thomas Wagner, Bonpote, 2021) can harm the reputation of a company involved in the practice. These factors make greenwashing a risk factor that can affect investments and that should be monitored at the company and portfolio levels.
The question thus arises to what extent greenwashing can be detected automatically. Such automated detection could help investors, journalists, prosecutors, and activists spot potentially illicit behavior, but it could also help companies avoid unintentionally misleading communication. In this paper, we provide a comprehensive survey of natural language processing (NLP) methods to identify potentially misleading climate-related corporate communication that can be indicative of greenwashing. We have identified 61 works that propose ways to check one or more aspects of greenwashing in text. We cover the intermediary steps involved in a natural language processing (NLP) analysis of the environmental disclosure of a company, ranging from detecting environmental or climate-related text to greenwashing detection. We structure these works in a unified categorization and discuss tasks, datasets, methods, and results for each of them. Unlike other surveys on the topic (Moodaley and Telukdarie, 2023b; Bernini et al., 2023), which compile and discuss findings reported in the scientific literature, our approach dives into the actual methodologies of these works.
Our main contributions and findings are:
  • We provide the first survey on NLP-based detection of greenwashing that describes and systemizes the methods used in the literature. Our survey focuses on tasks and datasets but also contextualizes the performance of the models and provides future research directions.
  • We find no datasets of positive and negative examples of greenwashing. Hence, it is not possible to train or evaluate models on the task of identifying greenwashing.
  • While there are no datasets explicitly labeled for greenwashing, many works have studied intermediary tasks, such as detecting misleading claims, vague language, or environmental reporting inconsistencies. Researchers have used these characteristics to develop theoretical definitions of greenwashing and proposed measurable indicators to identify potential cases. Although not a direct substitute for labeled greenwashing datasets, these indicators provide tools for analyzing corporate disclosures and assessing the likelihood of greenwashing.
  • Open challenges remain: While many models have been proposed for greenwashing-related tasks, it is difficult to compare approaches and solutions as many studies do not report measures of uncertainty (e.g., confidence interval, multiple runs, hyper-parameters tuning), nor do they report simple baselines that help assess the difficulty of datasets such as random baseline or TF-IDF baselines. Even the intermediary tasks thus remain active areas of research.
Our survey is structured as follows. We start with a discussion of the related work on greenwashing surveys (Section 2), followed by the legal context around climate actions and the definition of greenwashing (Section 3). We then survey intermediate tasks that have been studied in the literature (Section 4): identifying climate-related statements; identifying topics aligned with frameworks such as the Task Force on Climate-related Financial Disclosures (TCFD) or Environmental, Social, and Governance (ESG) factors; assessing the presence of green claims; and examining sentiment, argumentation quality, deceptive practices, and stance to discern the tone and authenticity of a message. We then review approaches that aim to identify greenwashing as a whole (Section 5). We conclude by discussing challenges and opportunities in greenwashing detection (Sections 6 and 7). We note that we have an Appendix section that will be part of the supplemental online-only material.
2. Related Work
@@
Moodaley and Telukdarie (2023b) analyze the literature on greenwashing, sustainability reporting, and the intersection with research in Artificial Intelligence (AI) and Machine Learning (ML). The authors provide an overview of how the literature has changed over time by examining the number of publications and the different keywords covered by such publications. Unlike our work, this survey does not describe the actual approaches of the publications. Bernini et al. (2023) review the definitions of greenwashing in business, management, and accounting. They conclude that there is no agreed-upon definition of greenwashing, except that it is defined as a gap between the company’s actions and its disclosure. In contrast to our study, the review provides only a brief overview of the techniques for detecting greenwashing. While our review focuses on the semantic and linguistic aspects of greenwashing identification, other works focus on quantitative approaches that can serve the same objective (Dao et al., 2024).
@@
Other works are less directly related to our survey. For example, a growing community is focused on using AI to tackle climate change [^1] See for example https://www.climatechange.ai/. Rolnick et al. (2022) review how machine learning has been used and could be used to tackle climate change. They explore different sectors of the economy, such as electricity, transportation, industry, and agriculture, among others. The survey discusses climate investments and the effects of climate change on finance, but not greenwashing. Alonso et al. (2023) provide a survey on ML approaches for climate finance, also known in the literature as green finance or carbon finance. Climate finance is a subfield in economics dedicated to providing tools from financial economics to tackle climate change. Finally, several works study the impact of AI on climate change (Hershcovich et al., 2022; Kaack et al., 2022; Rohde et al., 2023; Verdecchia et al., 2023). Our work, in contrast, surveys approaches that use AI to detect greenwashing.
3. Preliminaries
3.1. Legal Context
@@
The first international summit on the effect of humans on the environment was the United Nations Conference on the Human Environment, held in 1972 in Sweden (United Nations, European Commission, 2024). The Intergovernmental Panel on Climate Change (IPCC) was created in 1988 and released its first report in 1990  (Change, 1990). It was only in 1992 that the first treaty between countries was signed, the United Nations Framework Convention on Climate Change (UNFCCC) (United Nations, European Commission, 2024). The treaty only encouraged countries to reduce their emissions, and in 1997, the Kyoto Protocol set commitments for the countries to follow (United Nations Framework Convention on Climate Change, 2011). The 36 countries that participated in the Protocol reduced their emissions in 2008–2012 by a large margin in respect to the levels in 1990; however, the global emissions increased by $32\%$ in the same period. Since the creation of UNFCCC, the nations, i.e., the parties, have met yearly at the “Conference of the Parties” (COP).
The three main directions of climate change policy are reducing greenhouse emissions, promoting renewable energy, and improving energy efficiency. One major milestone was the introduction in 2005 of the European Union Emissions Trading System, the world’s first large-scale emissions trading scheme. The trading system limits the amount of CO2 emitted by European industries and covers $46\%$ of the EU’s CO2 emissions. It is estimated that the trading system reduced CO2 emissions in the EU by $3.8\%$ between 2008 and 2016 (Bayer and Aklin, 2020). At COP21, in 2015, 194 nations plus the European Union signed the Paris Agreement, a treaty by which the nations commit to keep the rise in global surface temperature below 2 °C (3.6 °F) above pre-industrial levels. To achieve this goal, each country sends a national climate action plan every five years, and the parties assess the collective progress made towards achieving the climate goals. The first such assessment took place at COP28, where it was established that an important direction to achieve the long-term goals of countries was to transition away from fossil fuels to renewable energy. In 2021, the European Climate Law was adopted: the EU commits to reducing its emissions by at least $55\%$ by 2030 with respect to the 1990 levels and becoming climate neutral by 2050 (Council of the EU Press release, 2021). Similarly, countries such as Canada (Government of Canada, 2024), Taiwan (The Republic of China Ministry of Environment(2023), Taiwan), South Korea (The Korean Law Information Center, 2024), and Australia (Government of New South Wales, 2023), among others, aim to achieve carbon neutrality by 2050 and have passed laws setting this goal.
@@
One of the first organizations to provide standards for reporting on climate change or social aspects is the Global Reporting Initiative (GRI), with its first guidelines published in 2000 (GRI, 2023). In 2015, the Financial Stability Board and the Group of 20 created the Task Force for Climate-related Financial Disclosure (TCFD) guidelines on disclosure in response to shortcomings of COP21, in particular the lack of standards climate-related disclosure (Wikipedia contributors, 2024). In 2021, the International Sustainability Standards Board (ISSB) was created to establish standards for climate-related disclosure, and starting in 2024, the standard released by this board will be applied worldwide (IFRS Foundation, 2024). ISSB standards were aligned with GRI disclosure standards to make them complementary and interoperable (Global Reporting Initiative, 2023), while the ISSB standards are expected to take over TCFD (IFRS Foundation, 2024). Differently from the voluntary reporting standards of GRI, TCFD, and ISSB, the Corporate sustainability reporting law in the EU required the creation of the European Sustainability Reporting Standards (ESRS), which are mandatory for companies subject to EU law (Directorate-General for Financial Stability, Financial Services and Capital Markets Union, 2023). ESRS has high interoperability with GRI and ISSB. GRI, TCFD, ESRS and ISSB fall under the umbrella term of environmental, social, and governance (ESG) guidelines for company disclosure. The European Union created its first law obliging companies to provide non-financial disclosure reporting in 2014, the Non-Financial Reporting Directive, which focused disclosure on environmental and social aspects. In 2023, the EU expanded this legislation via the Corporate Sustainability Reporting Law. Some European countries anticipated this legislation with their own, for example, the 2001 New Economic Regulations Act in France. Switzerland, which is not part of the EU, has imposed a mandatory TCFD disclosure for large public companies, banks, and insurance companies starting from January 2024 (The Swiss Federal Council, 2022), and similar laws exist in New Zealand (New Zealand Ministry of Environment, 2023). While not compulsory, initiatives like the Carbon Disclosure Project (CDP) have played a pivotal role by standardizing responses related to the climate disclosure of a company via structured questionnaires, thereby facilitating more systematic and comparable reporting.
3.2. Definition of Greenwashing
A widely cited and comprehensive definition, synthesizing those commonly found in the literature, is provided by the Oxford English Dictionary (OED, 2023).

### v2

# Detecting Greenwashing: A Natural Language Processing Literature Survey

@@
Abstract
Greenwashing refers to practices by corporations or governments that intentionally mislead the public about their environmental impact. This paper provides a comprehensive and methodologically grounded survey of natural language processing (NLP) approaches for detecting greenwashing in textual data, with a focus on corporate climate communication.
Rather than treating greenwashing as a single, monolithic task, we examine the set of NLP problems, also known as climate NLP tasks, that researchers have used to approximate it, ranging from climate topic detection to the identification of deceptive communication patterns. Our focus is on the methodological foundations of these approaches: how tasks are formulated, how datasets are constructed, and how model evaluation influences reliability.
Our review reveals a fragmented landscape: several subtasks now exhibit near-perfect performance under controlled settings, yet tasks involving ambiguity, subjectivity, or reasoning remain challenging. Crucially, no dataset of verified greenwashing cases currently exist.
We argue that advancing automated greenwashing detection requires principled NLP methodologies that combine reliable data annotations with interpretable model design. Future work should leverage third-party judgments, such as verified media reports or regulatory records, to mitigate annotation subjectivity and legal risk, and adopt decomposed pipelines that support human oversight, traceable reasoning, and efficient model design.
keywords:
Greenwashing detection, Natural Language Processing, Climate NLP, Climate communication, Corporate sustainability reporting
1 Introduction
@@
Since the first report of the Intergovernmental Panel on Climate Change in 1990 [change1990ipcc], world leaders and citizens have increasingly recognized the significant adverse impacts of human activities on the environment, particularly on climate change [climatebriefhistory]. This increased awareness has translated into laws and investments, such as the European Green Deal [europeangreendeal] or the Inflation Reduction Act in the US [inflationact]. Many companies have used the financial incentives offered by states, and the guidelines and legislation, to make significant steps towards sustainability [greencompanies]. At the same time, this growing attention also generated an advertising opportunity for companies that aim to promote themselves as environmentally aware and responsible. Indeed, some companies have been found to deliberately manipulate their data and statistics to appear more environment-friendly. The Diesel Scandal around the Volkswagen car company is a prominent example [volkswagen]. However, such cases are not the norm. More commonly, companies avoid outright data manipulation but present themselves in a misleadingly positive light regarding their environmental impact, a practice called greenwashing.
@@
Greenwashing is a widely observed practice rather than a marginal occurrence: the Australian Competition & Consumer Commission performed a sweep analysis on 247 businesses in October 2022 and found that $57\%$ of them made misleading claims about their environmental performance [sweepaustralia]. The International Consumer Protection Enforcement Network, too, found in November 2020 that $40\%$ of around 500 websites of businesses covering different sectors contain misleading green claims [globalsweep].
@@
Addressing greenwashing is not only in the broader public interest but also in the company’s best interest, as the practice is often subject to legal sanctions prohibiting misleading or deceptive conduct, with businesses facing fines or lawsuits as a consequence [litigatorsgreenwashing]. In addition, political awareness and public awareness of the issue via traditional and social media [bonpote] can harm the reputation of a company involved in the practice. These factors make greenwashing a risk factor that can affect investments and that should be monitored at the company and portfolio levels.
The growing public and regulatory attention to corporate climate claims raises a central question: to what extent can greenwashing be detected automatically? Automated detection systems could support investors, journalists, prosecutors, and activists in identifying potentially misleading or illicit communication, while also helping companies avoid unintentionally deceptive disclosures. Yet despite the relevance of this question, the methodological foundations of NLP-based greenwashing detection remain fragmented.
Methodological grounding
In this paper, we provide the first comprehensive and methodologically grounded survey of NLP techniques for identifying potentially misleading climate-related corporate communication. We analyze 61 scientific works that examine one or more aspects of greenwashing, ranging from identifying climate-related statements and assessing environmental claims to evaluating specificity, stance, argumentation, and the presence of deceptive linguistic cues. Unlike prior surveys [moodaley_greenwashing_2023, measuring_greenwashing], which primarily summarize findings reported in the literature, our focus is on the methodological foundations of these approaches: how tasks are formulated, how datasets are constructed, and how model evaluation influences reliability.
Summary of findings
Recent advances in NLP have enabled increasingly sophisticated analyses of climate-related corporate communication, yet our survey shows that progress across the building blocks of greenwashing detection is highly uneven. Several tasks, such as climate topic identification, climate risk classification, and the characterization of environmental claim types, now achieve near perfect performance when tackled with fine-tuned models. Even simple keyword-based approaches often achieve surprisingly strong results, suggesting that many of these tasks are driven by distinctive lexical terms rather than deep domain knowledge.
In contrast, tasks that require nuance, subjectivity, or reasoning, such as distinguishing specific from vague commitments or assessing rhetorical deception, remain challenging. Low inter-annotator agreement, ambiguous label definitions, and small datasets, undermine the reliability of the reported performance. This pattern reflects a broader issue: many real indicators of greenwashing require judgment on subtle differences, multi-document context, or implicit reasoning, where current models struggle and where evaluation is especially difficult.
These challenges are rendered more difficult by methodological weaknesses in research papers that limit the interpretability and reproducibility of results. Evaluation practices are often inadequate: few studies report uncertainty estimates, many rely on accuracy without considering semantic proximity between labels, and baselines are frequently missing. Only a minority compare to random, majority, or keyword-based baselines, even though these can drastically change how model performance is interpreted. Furthermore, many experiments are conducted on curated, balanced, and noise-free datasets, conditions that that might not hold in real-world datasets.
A more fundamental limitation is the absence of datasets containing verified instances of greenwashing. While the literature proposes theoretical indicators such as overly positive sentiment, rhetorical cues, and stance inconsistencies, none are validated against real cases. As a result, most studies evaluate proxy tasks rather than greenwashing, leaving the field without empirical foundations for end-to-end detection.
We believe that creating a greenwashing dataset requires principled data collection grounded in third-party judgments (e.g., regulators, courts, investigative journalism), clearer and narrower definitions of misleading communication, and decomposed pipelines that combine automated analysis with human oversight.
In conclusion, these findings show that while some subtasks appear solved, building a reliable and trustworthy NLP system for detecting greenwashing remains an open challenge.
Structure of the survey
The remainder of the paper proceeds as follows. We begin by situating our work within the broader literature on greenwashing and automated climate communication analysis (Section 2). We then outline the regulatory landscape that shapes how greenwashing is defined (Section 3), as these definitions constrain what can be annotated or modeled. Section 4 introduces the building-block tasks that underpin automated greenwashing detection: identifying climate-related statements; identifying topics aligned with frameworks such as the Task Force on Climate-related Financial Disclosures (TCFD) or Environmental, Social, and Governance (ESG) factors; assessing the presence of green claims; and examining sentiment, argumentation quality, deceptive practices, and stance to discern the tone and authenticity of a message. For each task, we give the task description, annotated or annotation-free datasets, models and an analysis of their performance, and finally insights, in which we highlight important limitations and potential future direction. Section 5 reviews approaches that aim to identify greenwashing as a whole. Finally, Sections 6 and 7 discuss open methodological challenges and outline research opportunities for building principled, reliable systems for greenwashing detecting.
2 Related Work
@@
moodaley_greenwashing_2023 analyze the literature on greenwashing, sustainability reporting, and the intersection with research in Artificial Intelligence (AI) and Machine Learning (ML). The authors provide an overview of how the literature has changed over time by examining the number of publications and the different keywords covered by such publications. Unlike our work, this survey does not describe the actual approaches of the publications. measuring_greenwashing review the definitions of greenwashing in business, management, and accounting. They conclude that there is no agreed-upon definition of greenwashing, except that it is defined as a gap between the company’s actions and its disclosure. In contrast to our study, the review provides only a brief overview of the techniques for detecting greenwashing. While our review focuses on the semantic and linguistic aspects of greenwashing identification, other works focus on quantitative approaches that can serve the same objective dao2024introduction. LUBLOY2025123399 studied the quantification of firm-level greenwashing; however, our work is focused on natural language processing approaches, while theirs is more generic.
@@
Other works are less directly related to our survey. For example, a growing community is focused on using AI to tackle climate change [^1] See for example https://www.climatechange.ai/. Rolnick_tackling_climate review how machine learning has been used and could be used to tackle climate change. They explore different sectors of the economy, such as electricity, transportation, industry, and agriculture, among others. The survey discusses climate investments and the effects of climate change on finance, but not greenwashing. alonsoMachineLearningMethods2023 provide a survey on ML approaches for climate finance, also known in the literature as green finance or carbon finance. Climate finance is a subfield in economics dedicated to providing tools from financial economics to tackle climate change. Finally, several works study the impact of AI on climate change [hershcovich_towards_2022, kaack_aligning_2022, rohde_broadening_2023, verdecchia_systematic_2023]. Our work, in contrast, surveys approaches that use AI to detect greenwashing.
3 Preliminaries
3.1 Legal Context
@@
The first international summit on the effect of humans on the environment was the United Nations Conference on the Human Environment, held in 1972 in Sweden [timelineclimate]. The Intergovernmental Panel on Climate Change (IPCC) was created in 1988 and released its first report in 1990 [change1990ipcc]. It was only in 1992 that the first treaty between countries was signed, the United Nations Framework Convention on Climate Change (UNFCCC) [timelineclimate]. The treaty only encouraged countries to reduce their emissions, and in 1997, the Kyoto Protocol set commitments for the countries to follow [kyotoprotocol]. The 36 countries that participated in the Protocol reduced their emissions in 2008–2012 by a large margin in respect to the levels in 1990; however, the global emissions increased by $32\%$ in the same period. Since the creation of UNFCCC, the nations, i.e., the parties, have met yearly at the “Conference of the Parties” (COP).
The three main directions of climate change policy are reducing greenhouse emissions, promoting renewable energy, and improving energy efficiency. One major milestone was the introduction in 2005 of the European Union Emissions Trading System, the world’s first large-scale emissions trading scheme. The trading system limits the amount of CO2 emitted by European industries and covers $46\%$ of the EU’s CO2 emissions. It is estimated that the trading system reduced CO2 emissions in the EU by $3.8\%$ between 2008 and 2016 [bayer2020european]. At COP21, in 2015, 194 nations plus the European Union signed the Paris Agreement, a treaty by which the nations commit to keep the rise in global surface temperature below 2 °C (3.6 °F) above pre-industrial levels. To achieve this goal, each country sends a national climate action plan every five years, and the parties assess the collective progress made towards achieving the climate goals. The first such assessment took place at COP28, where it was established that an important direction to achieve the long-term goals of countries was to transition away from fossil fuels to renewable energy. In 2021, the European Climate Law was adopted: the EU commits to reducing its emissions by at least $55\%$ by 2030 with respect to the 1990 levels and becoming climate neutral by 2050 [euclimatelaw]. Similarly, countries such as Canada [canadaclimatelaw], Taiwan [taiwanclimatelaw], South Korea [koreaclimatelaw], and Australia [australiaclimatelaw], among others, aim to achieve carbon neutrality by 2050 and have passed laws setting this goal.
@@
One of the first organizations to provide standards for reporting on climate change or social aspects is the Global Reporting Initiative (GRI), with its first guidelines published in 2000 [grihistory]. In 2015, the Financial Stability Board and the Group of 20 created the Task Force for Climate-related Financial Disclosure (TCFD) guidelines on disclosure in response to shortcomings of COP21, in particular the lack of standards climate-related disclosure [enwiki:1257716178]. In 2021, the International Sustainability Standards Board (ISSB) was created to establish standards for climate-related disclosure, and starting in 2024, the standard released by this board will be applied worldwide [ifrs]. ISSB standards were aligned with GRI disclosure standards to make them complementary and interoperable [gri], while the ISSB standards are expected to take over TCFD [ifrs]. Differently from the voluntary reporting standards of GRI, TCFD, and ISSB, the Corporate sustainability reporting law in the EU required the creation of the European Sustainability Reporting Standards (ESRS), which are mandatory for companies subject to EU law [esrs]. ESRS has high interoperability with GRI and ISSB. GRI, TCFD, ESRS and ISSB fall under the umbrella term of environmental, social, and governance (ESG) guidelines for company disclosure. The European Union created its first law obliging companies to provide non-financial disclosure reporting in 2014, the Non-Financial Reporting Directive, which focused disclosure on environmental and social aspects. In 2023, the EU expanded this legislation via the Corporate Sustainability Reporting Law. Some European countries anticipated this legislation with their own, for example, the 2001 New Economic Regulations Act in France. Switzerland, which is not part of the EU, has imposed a mandatory TCFD disclosure for large public companies, banks, and insurance companies starting from January 2024 [disclosureswiss], and similar laws exist in New Zealand [disclosurezealand]. While not compulsory, initiatives like the Carbon Disclosure Project (CDP) have played a pivotal role by standardizing responses related to the climate disclosure of a company via structured questionnaires, thereby facilitating more systematic and comparable reporting.
3.2 Definition of Greenwashing
A widely cited and comprehensive definition, synthesizing those commonly found in the literature, is provided by the Oxford English Dictionary [GreenwashMeaningsEtymology2023].